The data below pertain to two types of products manufactured by Cobb Corp. Fixed costs total $300,000 annually. The expected mix in units is 60% for product Y and 40% for product Z.

Per Unit

Sales Price

Variable Costs

Product Y

$120

$ 70

Product Z

500

200

How much is Cobb’s breakeven point in units?

A) 2,000

B)2,459

C) 857

D) 1,111
Answer #1
Y Z
Contribution margin(Sales price-Variable cost) (120-70)=$50 (500-200)=$300

Total Contribution margin=Respective Contribution margin*Respective sales mix

=(50*0.6)+(300*0.4)=$150

Hence breakeven point=Fixed cost/Total Contribution margin

=(300,000/150)

=2000